The length of an equilateral hexagon is 4 inches less than the the length of a side of a square. The perimeter of the hexagon is equal to the perimeter of the square. Find the length of a side of the hexagon and the length of a side of the square.
Answer by Earlsdon(6294)

Let the length of the side of the square = x, then the length of the side of the equilateral (regular) hexagon (six-sided polygon) is (x-4), so...
For the square, the perimeter (P) is:
P = 4x.
For the regular hexagon, the perimeter (P) is:
P = 6(x-4). Set the two perimeters equal.
6(x-4) = 4x Simplify and solve for x.
6x-24 = 4x
2x = 24
x = 12
The length of a side of the square = 12 inches.
The length of a side of the regular hexagon = x-4 = 12-4 = 8 inches.